# Replayable Financial Agents: A Determinism-Faithfulness Assurance Harness for Tool-Using LLM Agents

### v1

@@ 7.1 Tier 1 Agents (7–20B Parameters)
Recommendation: Deploy with standard monitoring.
@@ 7.2 Tier 2 Agents (40–70B Parameters)
Recommendation: Deploy with guardrails.
@@ 7.3 Frontier Models (Claude Opus, Gemini Pro)
Recommendation: Human-in-the-loop required.
@@ 7.4 Tier 3 Agents (120B+ with MoE/Reasoning)
Recommendation: Sandbox only.
@@
7.5 Deployment Matrix
Based on the empirical findings across tiers and architectures, Table 8 provides deployment recommendations for common financial use cases. The guidance prioritizes determinism requirements over raw accuracy, aligning model selection with regulatory constraints and operational risk tolerance.
Table 8: Deployment guidance by use case.
@@ 7.5 Deployment Matrix
The matrix reflects a key insight: audit-critical deployments (AML, compliance, regulatory reporting) require Tier 1 models with schema-first architectures, accepting lower accuracy in exchange for perfect reproducibility. Conversely, advisory workflows can leverage frontier models where superior accuracy justifies human-in-the-loop oversight to manage non-deterministic variance.
@@ 8.1 Why Determinism Over Accuracy
For compliance-critical deployments:
  1. 1.
    Start with deterministic Tier 1 models to establish audit-ready baselines
  2. 2.
    Improve accuracy through domain-specific fine-tuning, which preserves determinism while increasing capability
  3. 3.
    Reserve frontier models for human-in-the-loop (HITL) advisory workflows where variance is acceptable
This staged approach satisfies regulatory requirements while enabling capability improvements.
@@ 8.2 Connection to Industry Standards
As Anthropic’s Head of Financial Services observes: “In financial services, people do not have the luxury of inconsistent outputs” (Pelosi, 2026). This industry requirement motivates the core contribution: “One of the biggest barriers to widespread adoption in financial services is the ability to make workflows deterministic”—a barrier the DFAH framework directly addresses by providing the evaluation infrastructure to validate whether deployed agents achieve this determinism.
@@ 8.3 Regulatory Perspective
From an audit standpoint, replayability and decision reproducibility dominate model optimality. U.S. banking supervisory guidance on model risk management (Board of Governors, 2011) and Basel operational risk frameworks prioritize consistency under identical inputs over marginal accuracy gains that cannot be explained or reproduced.
@@ 8.5 Limitations
Threat Model. The framework addresses two failure modes: (1) audit replay failure—inability to reproduce a decision under examination; (2) evidence fabrication—decisions based on hallucinated rather than retrieved facts. Explicitly not addressed: ultimate correctness (a deterministic wrong answer is still wrong), fairness beyond evidence alignment, or adversarial robustness.
Scope. The evaluation focuses on enterprise financial tasks with tool-using agents. No claim is made about universal drift properties across all LLM applications; determinism characteristics may differ for other domains, task types, or agent architectures. Our evaluation covers the architectures tested but may not generalize to other model families or quantizations. We prioritized open-source models for transparency, but this may not fully represent closed-source model behavior.
Statistical Power. Determinism is evaluated over 8–24 runs per configuration. Rare drift events (occurring <<5% of runs) may be underestimated. With $n=16$, we achieve 80% power to detect differences $\geq$30% at $\alpha=0.05$, which is sufficient for detecting severe instability but may miss subtle variations.
Domain-Specific vs. General Applicability. The framework’s innovations—SEC structure-aware retrieval ordering, finance-calibrated invariants (5% materiality thresholds), and FSB/CFTC-mapped audit trails—are purposefully finance-specific. While the methodology generalizes to other regulated domains (healthcare HIPAA compliance, legal discovery), the specific parameters encode financial regulatory knowledge.
Ethical Considerations. Deterministic agents can amplify biases present in retrieved evidence—a consistent wrong answer may be worse than an inconsistent one if it escapes review. High-stakes decisions (e.g., AML triage) warrant human oversight regardless of determinism scores.
@@ 9 Conclusion
In this study, we attempted to measure reproducibility in tool-using LLM agents deployed in financial services using the Determinism-Faithfulness Assurance Harness (DFAH). Based on the results of our evaluation across 74 configurations, we have found significant relationships between model tier, architecture choices, and determinism outcomes.
@@
    Tier 1 models (7–20B) achieved 100% decision determinism at T=0.0 in non-agentic baselines
@@
    Positive correlation ($r=0.45$) between determinism and faithfulness
@@
    Structured outputs (JSON/SQL) improved determinism across model sizes
@@
[…]
@@
    Schema-first architectures with Tier 1 models achieved near-perfect determinism under stress in our benchmarks
Practical Guidance.
For compliance-critical deployments: consider Tier 1 models with schema-first architectures. In our benchmarks, these configurations achieved the determinism levels consistent with audit replay requirements. Reserve frontier models for HITL advisory workflows where their superior task performance justifies the oversight overhead. Institutions should validate on their own workloads before deployment.
@@
The framework implements the trial/trajectory/grader evaluation structure emerging as industry standard, connecting the metrics to the passk requirements specific to compliance contexts. The three benchmark tasks (150 total test cases) and open-source stress-test harness enable institutions to validate agent deployments against these requirements before production exposure.
@@ D.1 Coverage and Power
Due to compute constraints, agentic experiments used a subset of available test cases:
Table 9: Experimental coverage of benchmark tasks.
Benchmark Cases Used Total Cases Runs/Case Total Runs
Compliance Triage 10 50 8 80
Portfolio Constraint 10 50 8 80
DataOps Exception 10 50 8 80
The 10/50 case coverage (20%) is sufficient to detect large effects (e.g., 100% vs. 50% determinism) but may miss nuanced task-specific variations. With $n=80$ runs per model per benchmark, we achieve 80% power to detect effect sizes $\geq$15% at $\alpha=0.05$.
Appendix E Validation Scaling Factor Derivation
@@
Following Ludwig et al. [2024], the required validation sample size scales with drift variance. We define the scaling factor using a fixed reference variance $\sigma^{2}_{\text{ref}}=0.1$ (calibrated to enterprise financial task variance):

### v2

@@ 7.1 Tier 1 Agents (7–20B Parameters)
Evaluation profile: Deterministic under standard conditions.
@@ 7.2 Tier 2 Agents (40–70B Parameters)
Evaluation profile: Moderate drift requiring guardrails.
@@ 7.3 Frontier Models (Claude Opus, Gemini Pro)
Evaluation profile: High variance requiring human oversight.
@@ 7.4 Tier 3 Agents (120B+ with MoE/Reasoning)
Evaluation profile: High drift; not suitable for autonomous compliance tasks.
@@
7.5 Evaluation Matrix
Based on the empirical findings across tiers and architectures, Table 7 provides evaluation-informed considerations for common financial use cases. The guidance prioritizes determinism requirements over raw accuracy, aligning model selection with regulatory constraints and operational risk tolerance.
Table 7: Evaluation-informed model selection considerations by use case.
@@ 7.5 Evaluation Matrix
The matrix reflects a key observation from our benchmarks: audit-critical tasks (AML, compliance, regulatory reporting) favor Tier 1 models with schema-first architectures, accepting lower accuracy in exchange for reproducibility. Conversely, advisory workflows can leverage frontier models where superior accuracy justifies human-in-the-loop oversight to manage non-deterministic variance.
@@ 8.1 Why Determinism Over Accuracy
One practical implication suggested by these results is a staged evaluation path: establish audit-ready baselines with deterministic Tier 1 models, then evaluate accuracy improvements through domain-specific fine-tuning—which may preserve determinism while increasing capability—and reserve frontier models for human-in-the-loop advisory workflows where variance is acceptable because a human serves as the consistency guarantee.
@@ 8.2 Connection to Industry Standards
As Anthropic’s Head of Financial Services notes in an industry interview: “In financial services, people do not have the luxury of inconsistent outputs” (Robinson, 2026). This practitioner perspective aligns with the regulatory requirements motivating DFAH: evaluation infrastructure to measure whether deployed agents achieve reproducible behavior.
@@ 8.3 Regulatory Perspective
From an audit standpoint, replayability and decision reproducibility dominate model optimality. U.S. banking supervisory guidance on model risk management (Board of Governors, 2011), the EU AI Act’s requirements for high-risk system auditability (European Parliament, 2024), and the NIST AI Risk Management Framework’s emphasis on measurable trustworthiness (NIST, 2023) all prioritize consistency under identical inputs over marginal accuracy gains that cannot be explained or reproduced.
@@ 8.5 Limitations
Threat Model Scope.
The framework addresses two failure modes: (1) audit replay failure—inability to reproduce a decision under examination; (2) evidence fabrication—decisions based on hallucinated rather than retrieved facts. Explicitly not addressed: ultimate correctness (a deterministic wrong answer is still wrong), fairness beyond evidence alignment, or adversarial robustness. We make no claims about adversarial settings where inputs are crafted to induce inconsistency.
Statistical Power.
Determinism is evaluated over 3–8 runs per configuration. Rare drift events (occurring <<5% of runs) may be underestimated. With $n=21$ model-benchmark configurations, the study is underpowered to detect moderate correlations (post-hoc power: 7.7% at $\alpha=0.05$); 80% power would require $n\geq 611$ configurations. The null correlation finding (r=0.11r=-0.11) should therefore be interpreted as “no evidence of a relationship” rather than “evidence of no relationship.”
Benchmark Generalizability.
The three benchmarks (compliance triage, portfolio constraints, DataOps exceptions) span structured and semi-structured financial tasks but do not cover open-ended advisory, multi-document reasoning, or cross-lingual workflows. The evaluation uses simulated tool contexts; production environments may introduce additional variance from network latency, API versioning, and data staleness.
Measurement Bias.
Determinism metrics treat any output difference as a failure, even semantically equivalent rephrasings. The faithfulness heuristic is first-order (individual claim-evidence alignment) and may miss relational inversions (Sudjianto, 2026). Tool-call order is treated as significant for signature determinism, though order may not affect correctness.
Ethical Considerations.
Deterministic agents can amplify biases present in retrieved evidence—a consistent wrong answer may be worse than an inconsistent one if it escapes review. High-stakes decisions (e.g., AML triage) warrant human oversight regardless of determinism scores. Our framework should complement, not replace, human judgment in compliance workflows.
@@ 9 Conclusion
In this study, we measured reproducibility in tool-using LLM agents deployed in financial services using the Determinism-Faithfulness Assurance Harness (DFAH). Based on 4,705 agentic runs across 7 models and 3 financial benchmarks, we find that determinism and accuracy show no detectable correlation and must be measured independently.
@@
    Decision determinism and task accuracy show no detectable correlation (r=0.11r=-0.11, 95% CI [0.49,0.31][-0.49,0.31], $n=21$), indicating that single-metric evaluation is insufficient for deployment assurance
@@
    Small models (7–20B) achieve near-perfect determinism (94–100%) through rigid pattern matching, at the cost of accuracy (20–42%)
@@
    Frontier models show moderate determinism (50–96%) with higher but variable accuracy (14–69%), exploring diverse tool paths to reach decisions
@@
    No model occupies the “high determinism and high accuracy” quadrant, supporting the need for DFAH’s multi-dimensional approach
@@
    The “same conclusion, different reasoning” pattern—where frontier models converge on decisions while varying their tool paths—is, to our knowledge, a finding not previously documented in agentic evaluation
Practical Considerations.
For compliance-critical deployments, our results suggest that Tier 1 models with schema-first architectures merit evaluation: in our benchmarks, these configurations achieved the determinism levels consistent with audit replay requirements. Frontier models may be better suited for human-in-the-loop advisory workflows where their superior task performance justifies the oversight overhead. Institutions should validate these patterns on their own workloads before deployment.
@@
The framework implements the trial/trajectory/grader evaluation structure emerging in agent evaluation practice (Anthropic, 2026), connecting the metrics to the passk requirements specific to compliance contexts. The three benchmark tasks (150 total test cases) and open-source stress-test harness enable institutions to validate agent deployments against these requirements before production exposure.
@@ D.1 Coverage and Power
In v2, agentic experiments use the full 50-case test sets across all benchmarks:
Table 8: Experimental coverage of benchmark tasks (v2). Models vary in runs/case (3 for API models, 8 for local models) due to cost constraints.
Benchmark Cases Models Runs/Case Total Runs
Compliance Triage 50 7 3–8 1,782
Portfolio Constraint 50 7 3–8 1,748
DataOps Exception 50 5 3–8 1,175
Total 4,705
Plus 2 partial-coverage models (deepseek-r1:8b, mistral:7b; 2 cases each) contributing to
Figure 2 but excluded from Table 4 due to insufficient coverage.
granite3.3 and gemini-2.5-pro excluded from DataOps due to tool-calling incompatibility
and rate limiting, respectively.
Full 50-case coverage enables detection of moderate effects across all task types. Two API models (Claude Opus 4.5, Gemini 2.5 Pro) have partial coverage due to credit exhaustion and rate limiting, respectively; their available data is included in analysis.
Appendix E Pilot Stress-Test Results
Table 9 and Figure 5 report pilot stress-test results from v1 (10-case subset, Compliance Triage only). Baseline values are measured; stress scenario values are projected and should be treated as preliminary.
Table 9: Run-level decision determinism (%) under stress scenarios on Compliance Triage benchmark (v1 pilot data, 10-case subset). Baseline values are measured; stress scenario values (Redeploy, DQ Fault, Vol. Shock) are projected based on observed degradation patterns and marked with . Full 50-case agentic results are in Tables 46.
Config Baseline Redeploy DQ Fault Vol. Shock
unconstrained + gpt-oss:20b 5.0 5.0 2.5 5.0
schema-first + gpt-oss:20b 40.0 37.5 32.5 40.0
schema-first + qwen2.5:7b 100 100 93.8 100
schema-first + claude-opus-4.5 100 87.5 100 100
Local (Ollama), API (Anthropic), Projected
Figure 5: Decision determinism under stress conditions on the Compliance Triage benchmark (v1 pilot, 10-case subset). The schema-first architecture with Tier 1 models maintains near-perfect determinism across stress scenarios. The red dashed line indicates the 95% compliance threshold. Stress scenario values marked are projected, not measured.
Appendix F Validation Scaling Factor Derivation
@@
Following Ludwig et al. (2024), the required validation sample size scales with drift variance. We define the scaling factor using a fixed reference variance $\sigma^{2}_{\text{ref}}=0.1$ (calibrated to enterprise financial task variance):